OpenStudy (anonymous):

Factor the expression, if possible. Factor out any GCF first (including −1 if the leading coefficient is negative). (If not possible, enter the original expression.) 5m^8 + 29m^4 − 42

OpenStudy (anonymous):

Note that \(5m^8+29m^4-42 = 5(m^4)^2+29(m^4)-42\). If we make a change of variables, say \(u=m^4\), then we end up with a quadratic expression \(5u^2+29u-42.\) Try to factor this quadratic first, if possible. Once you factor it, remember to resubstitute \(m^4\) for \(u\) and try to factor the remaining parts in terms of \(m\). Can you try to take things from here? :-)

OpenStudy (anonymous):

no @ChristopherToni

OpenStudy (fibonaccichick666):

@ChristopherToni is correct. Try it @ludvic